OpenStudy (anonymous):

Solve the equation for p. Identify any extraneous solutions. Square root p=-1 A)-1 is a solution of the original equation. 1 is an extraneous solution. B) 1 is a solution of the original equation. C) 1 is a solution of the original equation. -1 is an extraneous solution. D) There is no solution.

jhonyy9 (jhonyy9):

sqrt(p) = -1 squaring both sides so what will get ?

jhonyy9 (jhonyy9):

sqrt(p) = p^(1/2) writing the squareroot like exponent so than now if you squared both sides what will get ?

jhonyy9 (jhonyy9):

do you like cooperat anything please ?

jhonyy9 (jhonyy9):

(p^(1/2) )^2 = (-1)^2 can you calculi it now ?

OpenStudy (anonymous):

I was on phone. I am mother that doesnt get some things. Dont be so rude if you dont want to answer dont

jhonyy9 (jhonyy9):

ok (p^(1/2))^2 = p^(1/2 *2) = p^(2/2) =p^1 = p

jhonyy9 (jhonyy9):

and (-1)^2 = -1 *-1 =